find the value of x in the equation 2(x-3)+5x=5(2x+6).
Answer by rmromero(383)


find the value of x in the equation 2(x-3)+5x=5(2x+6).
Simplify the equation given:
2x - 6 + 5x = 10x + 30 combine like terms
7x - 6 = 10x + 30 subtract 7x and 30 both sides
-36 = 3x divide both sides by 3
-12 = x
Checking :
2(x-3)+5x=5(2x+6)

2((-12)-3)+5(-12)=5(2(-12)+6)
2(-15) + 5(-12) = 5(-24+6)
-30 - 60 = 5(-18)
-90 = -90 ------> true
therefore, x = -12

If this solution helps,
Visit my website
http://onlinemathtutorial.tripod.com
Thanks